$621.00Giuseppis Pizza had orders for 5621.00 of pizzas. The prices were 519 for a large piz28, 315 for a medium pizza, and $12 for a small pizza. The number of large pizzas was two less than four times the number of medium pizza9. The number of small pizzas was two more than three times the number of medium pzzas. How many of each size of pizza were ordered? The number of medium siza pizzas i Type a whole nuber.) The number of large pizzas is Type a whole number) Th.ต number of small pizzas is D Type a whole nurmber.)
Answer #1

Let number of Medium sized pizzas = x
Then large pizzas = 4x - 2
Then medium pizzas = 3x+2

Price of Pizzas = 19*(4x-2) + 15x + 12*(3x+2) = 621
76x -38 + 15x + 36x + 24 = 621
127x = 621+38-24
127x = 635
x = 635/127 = 5

Number of Large pizzas = 4x-2 = 4*5-2 = 18
Number of Medium pizzas = x = 5
Number of Small Pizzas = 3x+2 = 3*5+ 2 = 17

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well